4. [10pts Find a tight asymptotic upper bound for the function T(n) defined by the recurence relation T(2)-2 T(n) = T(n/2) + Tuv )) + n Assume that n is a power of 2
Answer #1

We expand the recursive formula and find the following

T(n) = n + n n + + upto log_2 n-1 terms +

                 \frac{\sqrt{n}}{2}+\frac{\sqrt{n}}{4}+\frac{\sqrt{n}}{8}+... upto log*_2 n-1 terms +

                 ... +

                n^{\frac{1}{2}}+n^{\frac{1}{4}}+... upto log_2 n-1 terms

           =n\times \frac{1-(\frac{1}{2})^{log_2 n}}{1-\frac{1}{2}}+\frac{\sqrt{n}}{2}\times \frac{1-(\frac{1}{2})^{log_2 n}}{1-\frac{1}{2}}+...+\sqrt{n}

            =2n+\sqrt{n}+...+\sqrt{n}

           =O(n+\sqrt{n}) by ignoring smaller terms